# PlainMamba: Improving Non-Hierarchical Mamba in Visual Recognition

### v1

@@ 3 Method
In this section, we introduce PlainMamba in detail. An overview of the proposed model is presented in Figure 2 (a). In Section 3.1, we revisit the background of State Space Modeling and Mamba. Then, we introduce our model architecture and block architecture in Section 3.2 and Section 3.3, respectively. Finally, in Section 3.4, we present different variants of PlainMamba.
@@ 3.1 Preliminaries
3.1.1 State Space Models.
SSMs are typically used to model a continuous linear time-invariant (LTI) system [90] where an input signal $x(t)\in\mathbb{R}$ is mapped to its output signal $y(t)\in\mathbb{R}$ through a state variable $h(t)\in\mathbb{R}^{m}$ with the following rules:
@@
where $\mathbf{A}\in\mathbb{R}^{m\times m}$, $\mathbf{B}\in\mathbb{R}^{m\times 1}$, $\mathbf{C}\in\mathbb{R}^{1\times m}$ and $\mathbf{D}\in\mathbb{R}^{1\times 1}$ are parameters.
To make the above system usable for a discrete system, e.g., a sequence-to-sequence task, a timescale parameter $\mathbf{\Delta}$ is used to transform the parameters $\mathbf{A}$ and $\mathbf{B}$ to their discretized counterparts $\bar{\mathbf{A}}$ and $\bar{\mathbf{B}}$. In Mamba [31] and its following works [58, 108], this is achieved with the following zero-order hold (ZOH) rule:
@@
3.1.2 Mamba.
Since SSMs are often used to model LTI systems, their model parameters are shared by all time steps $i$. However, as found in Mamba [31], such time-invariant characteristics severely limit the model’s representativity. To alleviate this problem, Mamba lifts the time-invariant constraint and makes the parameters $\mathbf{B}$, $\mathbf{C}$ and $\mathbf{\Delta}$ dependent on the input sequence $\{x_{i}\}$, a process they refer to as the selective scan, resulting in the token-dependent $\{\mathbf{B}_{i}\}$, $\{\mathbf{C}_{i}\}$ and $\{\mathbf{\Delta}_{i}\}$. Moreover, the SSM is combined with a gated MLP [43] to gain better representation ability. Specifically, the output sequence $\{y_{i}\}$ is computed from the $\{x_{i}\}$ as the following [^1] For simplicity, here we assume $x_{i}$ to have a single channel. It is straightforward to modify the formulation for the multi-channel case. The output linear projection that maps the features to their original dimension is also omitted here.:
@@
Figure 2: (a) The overall architecture of the proposed PlainMamba. PlainMamba does not have a hierarchical structure, it instead stacks $L$ identical PlainMamba block to form the main network. For image classification, it uses global average pooling instead of the CLS to gather global information. (b) Architecture of PlainMamba block, which is similar to the Mamba [31] block where the selective scanning is combined with a gated MLP. (c) The proposed Direction-Aware Updating, where a series of learnable parameters $\mathbf{\Theta}_{k}$ are combined with the data-dependent updating parameters to explicitly inject relative 2D positional information into the selective scanning process.
@@ 3.2 Overall architecture of PlainMamba
In Figure 2, we present the model architecture of PlainMamba. Our model is divided into three main components: (1) a convolutional tokenizer that transforms an input 2D image into a sequence of visual tokens, (2) the main network with a series of $L$ identical PlainMamba blocks to learn high-quality visual representations; and (3) a task-specific head for different types of downstream tasks.
In more detail, the tokenizer will downsample the input image $I\in\mathbb{R}^{H_{I}\times W_{I}\times 3}$ into a list of visual tokens $x\in\mathbb{R}^{H\times W\times C}$, where $C$ is the channel number. We set the default down-sampling factor to 16, following ViT [24]. After combining the initial visual tokens with positional embeddings [82] for retaining spatial information, the tokens undergo a series of transformations through the $L$ PlainMamba blocks, which are designed to simplify usage by maintaining the input-output shape consistency. The final stage of the architecture involves a task-specific head, which is dependent on the particular downstream application. For instance, in image classification tasks, the image tokens are globally pooled into a vector, which is then fed into a linear classification head to produce the final output.
PlainMamba distinguishes itself from existing vision transformers [24, 57] and concurrent vision Mamba [58, 108] architectures in several key aspects. Firstly, it does not use any special tokens, such as the commonly used CLS token. Secondly, in contrast to approaches that adopt a hierarchical structure to manage feature resolution [87, 51, 59], Instead, PlainMamba maintains a constant feature resolution across all blocks. This design choice considers the recent progress made in various visual foundation models [46, 65, 63] where the plain non-hierarchical ViT is used rather than its hierarchical counterparts.
@@ 3.3 PlainMamba Block
The overall PlainMamba architecture is comprised of several identical PlainMamba blocks, which form the backbone for learning high-quality visual features. We present the structure of the PlainMamba block in Figure 2. We take inspiration from the original Mamba block where the SSM is combined with a gated MLP in a unified module. However, we make several key adjustments to the selective scanning process to fully exploit the two-dimensional nature of image inputs. This adaptation is crucial for effectively transitioning from the inherently 1D processing paradigm of language models to the 2D domain of images. To this end, we introduce two novel techniques: (1) Continuous 2D Scanning and (2) Direction-Aware Updating. The first technique ensures that each visual token is always adjacent to the previous scanned token. By doing so, it mitigates positional bias and encourages a more uniform understanding of the image space, enhancing the model’s ability to learn from visual inputs. The second technique explicitly embeds the 2D relative positional information into the selective scanning process, which allows the model to better interpret the positional context of flattened visual tokens.
Figure 3: Comparison between our Continuous 2D Scanning and the selective scan orders in ViM [108] and VMamba [58]. Our method makes sure that every scanned visual token is spatially adjacent to its predecessor, avoiding potential spatial and semantic discontinuity.
3.3.1 Continuous 2D Scanning
The selective scan mechanism is inherently designed for sequential data, such as text. Adapting this mechanism for 2D image data requires flattening the 2D viusal tokens into a 1D sequence to apply the State Space Model (SSM) updating rule. Prior research, e.g., VisionMamba [108] and VMamba [58], has demonstrated the efficacy of using multiple scanning orders to enhance model performance — such as both row-wise and column-wise scans in multiple directions. However, as shown in Figure 3 (a) and (b), in these approaches, each scanning order can only cover one type of 2D direction, e.g., left to right, causing spatial discontinuity when moving to a new row (or column). Moreover, as the parameter $\mathbf{A}$ in Equation 3 serves as a decaying term, such spatial discontinuity can also cause adjacent tokens to be decayed to different degrees, compounding the semantic discontinuity and resulting in potential performance drop.
@@ 3.3.1 Continuous 2D Scanning
3.3.2 Direction-Aware Updating
As shown in Equation 3, the contribution of a token $x_{i}$ to the hidden state $h_{i}$ in the selective scan is determined by the parameter $\bar{\mathbf{B}}_{i}$, derived from $x_{i}$ itself. In language models, the sequential order naturally dictates the positional relationship between tokens, allowing the model to "remember" their relative positions. However, in our Continuous 2D Scanning, the current token can be in one of four possible directions relative to its predecessor. This challenges the model’s ability to discern the precise spatial relationship between consecutive tokens based on $\mathbf{B}_{i}$ alone.
Our Direction-Aware Updating is therefore proposed to address this challenge. Drawing inspiration from the relative positional encoding mechanisms in vision transformers [24], we employ a set of learnable parameters $\{\mathbf{\Theta}_{k}\in\mathbb{R}^{m\times 1}\}$ (for $k=1,2,...,5$), representing the four cardinal directions plus a special BEGIN direction for the initial token. These parameters are summed with the data-dependent $\mathbf{B}_{i}$ to enrich the selective scan process with directional information. Specifically, with $x_{i}$ and $z_{i}$ following Equation 3, our Direction-Aware Updating is formulated as follows:
@@ 3.3.2 Direction-Aware Updating
Model Hierarchical Params FLOPs Top-1 CNN ResNeXt101-32$\times$[96] 44M 8.0G 78.6 ResNeXt101-32$\times$[96] 88M 16.5G 79.3 RegNetY-4G [66] 21M 4.0G 80.0 RegNetY-8G [66] 39M 8.0G 81.7 ConvNeXt-T [60] 29M 4.5G 82.1 ConvNeXt-S [60] 50M 8.7G 83.1 Transformer DeiT-Tiny [80] 5M 1.3G 72.2 DeiT-Small [80] 22M 4.6G 79.9 DeiT-Base [80] 86M 16.8G 81.8 Swin-Tiny [59] 29M 4.5G 81.3 Swin-Small [59] 50M 8.7G 83.0 PVT-Tiny [86] 13M 2G 75.1 PVT-Small [86] 25M 3.8G 79.8 PVT-Medium [86] 44M 6.7G 81.2 Focal-Tiny [101] 29M 4.9G 82.2 Focal-Small [101] 51M 9.1G 83.5 State Space Modeling ViM-T [108] 7M - 76.1 ViM-S [108] 26M - 80.5 LocalViM-T [45] 8M 1.5G 76.2 LocalViM-S [45] 28M 4.8G 81.2 Mamba-ND-T [49] 24M - 79.2 Mamba-ND-S [49] 63M - 79.4 S4ND-ViT-B [62] 89M - 80.4 S4ND-ConvNeXt-T [62] 30M - 82.2 VMamba-T [58] 22M 5.6G 82.2 VMamba-S [58] 44M 11.2G 83.5 PlainMamba-L1 7M 3.0G 77.9 PlainMamba-L2 25M 8.1G 81.6 PlainMamba-L3 50M 14.4G 82.3
3.4 Model Variants of PlainMamba
We present three different model variants of PlainMamba. We report their architectural details in Table 1. Specifically, from PlainMamba-L1 to PlainMamba-L2, we scale the model width, i.e., feature channel numbers, and keep the model depth to 24. From PlainMamba-L2 to PlainMamba-L3, we scale both model width and depth. The FLOPs are measured using 224$\times$224 inputs, and we follow the official Mamba codebase to compute the FLOPs of the selective scan process.
4 Experiments
4.1 ImageNet-1K Classification
@@
Setting: We build our codebase following [100]. Specifically, for the ImageNet-1k experiments, we train all PlainMamba models for 300 epochs using AdamW optimizer. Following [59], we set the batch size to 2048, weight decay to 0.05, and the peak learning rate to 0.002. Cosine learning rate scheduling is used. For data augmentation, we used the commonly used recipe [80, 59, 23, 22], which includes Mixup [104], Cutmix [102], Random erasing [106] and Rand augment [15].
Results: In Table 2, we report the ImageNet-1K experiment results. We compare PlainMamba with three different kinds of visual feature extractors: CNNs, vision transformers, and SSMs. In addition, the comparison includes both hierarchical and non-hierarchical models. Firstly, when comparing with SSMs, our model is doing better than the recently proposed Vision Mamba [108] and Mamba-ND [49]. For example, PlainMamba-L2 achieves a 2.4% higher accuracy than Mamba-ND-T while they share a similar model size. These results validate PlainMamba’s effectiveness as a non-hierarchical SSM. Secondly, when compared with CNNs and transformers, our model achieves better performance than the non-hierarchical counterparts. For example, PlainMamba-L2 achieves 1.7% better accuracy with DeiT-Small. Moreover, PlainMamba also achieves similar performance when compared with hierarchical models. For example, when the model size is around 25M, our model achieves 0.3% better accuracy than Swin-Tiny, validating PlainMamba’s ability as a general feature extractor. On the other hand, the hierarchical VMamba [58], together with other hierarchical transformers, do achieve a better accuracy than ours. As we explained in Section 1, hierarchical models tend to perform better than non-hierarchical ones in visual recognition. As the main motivation of our work is to develop a simple Mamba architecture, a bit inferior ImageNet accuracy is acceptable.
Table 3: Mask R-CNN object detection and instance segmentation on MS COCO mini-val using 1$\times$ schedule. We use ViTAdapter [11] to compute multi-scale features. FLOPs are computed using input size 1280$\times$800.
Backbone Hierarchical Params FLOPs $AP^{bb}$ $AP^{bb}_{50}$ $AP^{bb}_{75}$ $AP^{mk}$ $AP^{mk}_{50}$ $AP^{mk}_{75}$ CNN ResNeXt101-32x4d [96] 63M 340G 41.9 - - 37.5 - - ResNeXt101-64x4d [96] 102M 493G 42.8 - - 38.4 - - Transformer ViT-Adapter-T [11] 29M 349G 41.1 62.5 44.3 37.5 59.7 39.9 ViT-Adapter-S [11] 49M 463G 44.7 65.8 48.3 39.9 62.5 42.8 ViT-Adapter-B [11] 131M 838G 47.0 68.2 51.4 41.8 65.1 44.9 PVT-Small [86] 44M - 40.4 62.9 43.8 37.8 60.1 40.3 PVT-Medium [86] 64M - 42.0 64.4 45.6 39.0 61.6 42.1 PVT-Large [86] 81M - 42.9 65.0 46.6 39.5 61.9 42.5 Swin-Tiny [59] 48M 264G 42.2 - - 39.1 - - Swin-Small [59] 69M 354G 44.8 - - 40.9 - - ViL-Tiny [105] 26M 145G 41.4 63.5 45.0 38.1 60.3 40.8 ViL-Small [105] 45M 218G 44.9 67.1 49.3 41.0 64.2 44.1 ViL-Medium [105] 60M 293G 47.6 69.8 52.1 43.0 66.9 46.6 State Space Modeling EfficientVMamba-T [64] 11M 60G 35.6 57.7 38.0 33.2 54.4 35.1 EfficientVMamba-S [64] 31M 197G 39.3 61.8 42.6 36.7 58.9 39.2 EfficientVMamba-B [64] 53M 252G 43.7 66.2 47.9 40.2 63.3 42.9 VMamba-T [58] 42M 262G 46.5 68.5 50.7 42.1 65.5 45.3 VMamba-S [58] 64M 357G 48.2 69.7 52.5 43.0 66.6 46.4 PlainMamba-Adapter-L1 31M 388G 44.1 64.8 47.9 39.1 61.6 41.9 PlainMamba-Adapter-L2 53M 542G 46.0 66.9 50.1 40.6 63.8 43.6 PlainMamba-Adapter-L3 79M 696G 46.8 68.0 51.1 41.2 64.7 43.9
4.2 COCO object detection and instance segmentation

### v2

@@ 3.1 Preliminaries
State Space Models. SSMs are typically used to model a continuous linear time-invariant (LTI) system [Williams et al.(2007)Williams, Lawrence, et al.] where an input signal $x(t)\in\mathbb{R}$ is mapped to its output signal $y(t)\in\mathbb{R}$ through a state variable $h(t)\in\mathbb{R}^{m}$ with the following rules:
@@
where $\mathbf{A}\in\mathbb{R}^{m\times m}$, $\mathbf{B}\in\mathbb{R}^{m\times 1}$, $\mathbf{C}\in\mathbb{R}^{1\times m}$ and $\mathbf{D}\in\mathbb{R}^{1\times 1}$ are parameters. To make the above system usable for a discrete system, e.g., a sequence-to-sequence task, a timescale parameter $\mathbf{\Delta}$ is used to transform the parameters $\mathbf{A}$ and $\mathbf{B}$ to their discretized counterparts $\bar{\mathbf{A}}$ and $\bar{\mathbf{B}}$. In Mamba [Gu and Dao(2023)] and its following works [Liu et al.(2024c)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu, Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang], this is achieved with the following zero-order hold (ZOH) rule:
@@
Mamba. Since SSMs are often used to model LTI systems, their model parameters are shared by all time steps $i$. However, as found in Mamba [Gu and Dao(2023)], such time-invariant characteristics severely limit the model’s representativity. To alleviate this problem, Mamba lifts the time-invariant constraint and makes the parameters $\mathbf{B}$, $\mathbf{C}$ and $\mathbf{\Delta}$ dependent on the input sequence $\{x_{i}\}$, a process they refer to as the selective scan, resulting in the token-dependent $\{\mathbf{B}_{i}\}$, $\{\mathbf{C}_{i}\}$ and $\{\mathbf{\Delta}_{i}\}$. Moreover, the SSM is combined with a gated MLP [Hua et al.(2022)Hua, Dai, Liu, and Le] to gain better representation ability. Specifically, the output sequence $\{y_{i}\}$ is computed from the $\{x_{i}\}$ as the following:
@@ 3.2 Overall architecture of PlainMamba
In Figure 2, we present the model architecture of PlainMamba. Our model is divided into three main components: (1) a convolutional tokenizer that transforms an input 2D image into visual tokens, (2) the main network with a series of $L$ identical PlainMamba blocks to learn visual representations, and (3) a task-specific head for downstream applications.
In more detail, the tokenizer will downsample the input image $I\in\mathbb{R}^{H_{I}\times W_{I}\times 3}$ into a list of visual tokens $x\in\mathbb{R}^{H\times W\times C}$, where $C$ is the channel number. We set the default down-sampling factor to 16, following ViT [Dosovitskiy et al.(2021)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, Uszkoreit, and Houlsby]. After combining the initial visual tokens with positional embeddings [Vaswani et al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones, Gomez, Kaiser, and Polosukhin] for retaining spatial information, the tokens undergo a series of transformations through the $L$ PlainMamba blocks, which are designed to simplify usage by maintaining the input-output shape consistency. The final stage of the architecture involves a task-specific head, which is dependent on the particular downstream application. For instance, in image classification tasks, the image tokens are globally pooled into a vector, which is then fed into a linear classification head to produce the final output.
PlainMamba distinguishes itself from existing vision transformers [Dosovitskiy et al.(2021)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, Uszkoreit, and Houlsby, Liu et al.(2024a)Liu, Yang, Zhou, Xi, Yu, Yu, Liang, Shi, Zhang, Zheng, et al.] and concurrent vision Mamba [Liu et al.(2024c)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu, Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang] architectures in several key aspects. Firstly, it does not use any special tokens, such as the commonly used CLS token. Secondly, in contrast to approaches that adopt a hierarchical structure to manage feature resolution [Wang et al.(2022)Wang, Xie, Li, Fan, Song, Liang, Lu, Luo, and Shao, Li et al.(2022b)Li, Wu, Fan, Mangalam, Xiong, Malik, and Feichtenhofer, Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo], Instead, PlainMamba maintains a constant feature resolution across all blocks. This design choice considers the recent progress made in various visual foundation models [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, Dollár, and Girshick, Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al., Oquab et al.(2023)Oquab, Darcet, Moutakanni, Vo, Szafraniec, Khalidov, Fernandez, Haziza, Massa, El-Nouby, Howes, Huang, Xu, Sharma, Li, Galuba, Rabbat, Assran, Ballas, Synnaeve, Misra, Jegou, Mairal, Labatut, Joulin, and Bojanowski] where the plain non-hierarchical ViT is used rather than its hierarchical counterparts.
@@ 3.3 PlainMamba Block
The overall architecture comprises several identical PlainMamba blocks, forming the backbone for learning high-quality visual features. We present the structure of the PlainMamba block in Figure 2, in which we make several key adjustments to the original Mamba block to fully exploit the two-dimensional nature of image inputs. This adaptation is crucial for effectively transitioning from the inherently 1D processing paradigm of language models to the 2D domain of images. To this end, we introduce two novel techniques: (1) Continuous 2D Scanning and (2) Direction-Aware Updating. The first technique ensures that each visual token is always adjacent to the previous scanned token. By doing so, it mitigates positional bias and encourages a more uniform understanding of the image space, enhancing the model’s ability to learn from visual inputs. The second technique explicitly embeds the 2D relative positional information into the selective scanning process, which allows the model to better interpret the positional context of flattened visual tokens.
Figure 3: Comparison between our Continuous 2D Scanning and the selective scan orders in ViM [Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang] and VMamba [Liu et al.(2024c)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu]. Our method makes sure that every scanned visual token is spatially adjacent to its predecessor, avoiding potential spatial and semantic discontinuity.
Continuous 2D Scanning. The selective scan mechanism is inherently designed for sequential data, such as text. Adapting this mechanism for 2D image data requires flattening the 2D viusal tokens into a 1D sequence to apply the State Space Model (SSM) updating rule. Prior research, e.g., VisionMamba [Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang] and VMamba [Liu et al.(2024c)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu], has demonstrated the efficacy of using multiple scanning orders to enhance model performance — such as both row-wise and column-wise scans in multiple directions. However, as shown in Figure 3 (a) and (b), in these approaches, each scanning order can only cover one type of 2D direction, e.g., left to right, causing spatial discontinuity when moving to a new row (or column). Moreover, as the parameter $\mathbf{A}$ in Equation 3 serves as a decaying term, such spatial discontinuity can also cause adjacent tokens to be decayed to different degrees, compounding the semantic discontinuity and resulting in potential performance drop.
@@
Direction-Aware Updating. As shown in Equation 3, the contribution of a token $x_{i}$ to the hidden state $h_{i}$ in the selective scan is determined by the parameter $\bar{\mathbf{B}}_{i}$, derived from $x_{i}$ itself. In language models, the sequential order naturally dictates the positional relationship between tokens, allowing the model to "remember" their relative positions. However, in our Continuous 2D Scanning, the current token can be in one of four possible directions relative to its predecessor. This challenges the model’s ability to discern the precise spatial relationship between consecutive tokens based on $\mathbf{B}_{i}$ alone. Our Direction-Aware Updating is therefore proposed to address this challenge. Drawing inspiration from the relative positional encoding mechanisms in vision transformers [Dosovitskiy et al.(2021)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, Uszkoreit, and Houlsby], we employ a set of learnable parameters $\{\mathbf{\Theta}_{k}\in\mathbb{R}^{m\times 1}\}$ (for $k=1,2,...,5$), representing the four cardinal directions plus a special BEGIN direction for the initial token. These parameters are summed with the data-dependent $\mathbf{B}_{i}$ to enrich the selective scan process with directional information. Specifically, with $x_{i}$ and $z_{i}$ following Equation 3, our Direction-Aware Updating is formulated as follows:
@@
3.4 Model Variants of PlainMamba
As shown in Table 1, we present three different model variants of PlainMamba. Specifically, from PlainMamba-L1 to PlainMamba-L2, we scale the model width, i.e., feature channel numbers, and keep the model depth to 24. From PlainMamba-L2 to PlainMamba-L3, we scale both model width and depth. The FLOPs are measured using 224$\times$224 inputs, and we follow the official Mamba codebase to compute the FLOPs of the selective scan process.
4 Experiments
@@
Model Hierarchical Params FLOPs Top-1 CNN ResNeXt101-32$\times$[Xie et al.(2017)Xie, Girshick, Dollár, Tu, and He] 44M 8.0G 78.6 ResNeXt101-32$\times$[Xie et al.(2017)Xie, Girshick, Dollár, Tu, and He] 88M 16.5G 79.3 RegNetY-4G [Radosavovic et al.(2020)Radosavovic, Kosaraju, Girshick, He, and Dollár] 21M 4.0G 80.0 RegNetY-8G [Radosavovic et al.(2020)Radosavovic, Kosaraju, Girshick, He, and Dollár] 39M 8.0G 81.7 ConvNeXt-T [Liu et al.(2022)Liu, Mao, Wu, Feichtenhofer, Darrell, and Xie] 29M 4.5G 82.1 ConvNeXt-S [Liu et al.(2022)Liu, Mao, Wu, Feichtenhofer, Darrell, and Xie] 50M 8.7G 83.1 Transformer DeiT-Tiny [Touvron et al.(2021)Touvron, Cord, Douze, Massa, Sablayrolles, and Jégou] 5M 1.3G 72.2 DeiT-Small [Touvron et al.(2021)Touvron, Cord, Douze, Massa, Sablayrolles, and Jégou] 22M 4.6G 79.9 DeiT-Base [Touvron et al.(2021)Touvron, Cord, Douze, Massa, Sablayrolles, and Jégou] 86M 16.8G 81.8 Swin-Tiny [Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo] 29M 4.5G 81.3 Swin-Small [Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo] 50M 8.7G 83.0 PVT-Tiny [Wang et al.(2021)Wang, Xie, Li, Fan, Song, Liang, Lu, Luo, and Shao] 13M 2G 75.1 PVT-Small [Wang et al.(2021)Wang, Xie, Li, Fan, Song, Liang, Lu, Luo, and Shao] 25M 3.8G 79.8 PVT-Medium [Wang et al.(2021)Wang, Xie, Li, Fan, Song, Liang, Lu, Luo, and Shao] 44M 6.7G 81.2 Focal-Tiny [Yang et al.(2021)Yang, Li, Zhang, Dai, Xiao, Yuan, and Gao] 29M 4.9G 82.2 Focal-Small [Yang et al.(2021)Yang, Li, Zhang, Dai, Xiao, Yuan, and Gao] 51M 9.1G 83.5 State Space Modeling ViM-T [Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang] 7M - 76.1 ViM-S [Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang] 26M - 80.5 LocalViM-T [Huang et al.(2024)Huang, Pei, You, Wang, Qian, and Xu] 8M 1.5G 76.2 LocalViM-S [Huang et al.(2024)Huang, Pei, You, Wang, Qian, and Xu] 28M 4.8G 81.2 Mamba-ND-T [Li et al.(2024)Li, Singh, and Grover] 24M - 79.2 Mamba-ND-S [Li et al.(2024)Li, Singh, and Grover] 63M - 79.4 S4ND-ViT-B [Nguyen et al.(2022)Nguyen, Goel, Gu, Downs, Shah, Dao, Baccus, and Ré] 89M - 80.4 S4ND-ConvNeXt-T [Nguyen et al.(2022)Nguyen, Goel, Gu, Downs, Shah, Dao, Baccus, and Ré] 30M - 82.2 VMamba-T [Liu et al.(2024c)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu] 22M 5.6G 82.2 VMamba-S [Liu et al.(2024c)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu] 44M 11.2G 83.5 PlainMamba-L1 7M 3.0G 77.9 PlainMamba-L2 25M 8.1G 81.6 PlainMamba-L3 50M 14.4G 82.3
In the main paper, we quantitatively compare PlainMamba with previously proposed models on four visual recognition tasks: image classification, object detection, instance segmentation, and semantic segmentation. Please refer to our supplementary materials for further ablation studies.
4.1 Experiment Settings
@@
ImageNet Classification. We build our codebase following [Yang et al.(2023)Yang, Xu, Mello, Crowley, and Wang], which is a commonly used training recipe. Specifically, for the ImageNet-1k experiments, we train all PlainMamba models for 300 epochs using AdamW optimizer. Following [Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo], we set the batch size to 2048, weight decay to 0.05, and the peak learning rate to 0.002. Cosine learning rate scheduling is used. For data augmentation, we used the commonly used recipe [Touvron et al.(2021)Touvron, Cord, Douze, Massa, Sablayrolles, and Jégou, Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo, Dong et al.(2022)Dong, Bao, Chen, Zhang, Yu, Yuan, Chen, and Guo, Ding et al.(2022)Ding, Xiao, Codella, Luo, Wang, and Yuan], which includes Mixup [Zhang et al.(2017)Zhang, Cisse, Dauphin, and Lopez-Paz], Cutmix [Yun et al.(2019)Yun, Han, Oh, Chun, Choe, and Yoo], Random erasing [Zhong et al.(2020)Zhong, Zheng, Kang, Li, and Yang] and Rand augment [Cubuk et al.(2020)Cubuk, Zoph, Shlens, and Le].
ADE20K Semantic Segmentation. We follow common practice [Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo, Dong et al.(2022)Dong, Bao, Chen, Zhang, Yu, Yuan, Chen, and Guo, Yang et al.(2023)Yang, Xu, Mello, Crowley, and Wang] to use UperNet [Xiao et al.(2018)Xiao, Liu, Zhou, Jiang, and Sun] as the segmentation network. Unlike XCiT [Ali et al.(2021)Ali, Touvron, Caron, Bojanowski, Douze, Joulin, Laptev, Neverova, Synnaeve, Verbeek, et al.], we do not explicitly resize the constant resolution feature maps into multi-scale. Following [Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo], we train all models for 160 iterations with batch size 16 and set the default training image size to 512$\times$512.